OpenStudy (anonymous):

the perimeter of a rectangle is 68 ft. The ratio of its length to its width is 9:8. What are the dimensions of the rectangle? A.) 22 ft by 12 ft B.) 18 ft by 16 ft C.) 20 ft by 14 ft D.) 24 ft by 10 ft

OpenStudy (anonymous):

@kels200105 @aum @Agl202

OpenStudy (anonymous):

B duh

OpenStudy (anonymous):

its the only one that follows the ratio

OpenStudy (agl202):

18 ft by 16 ft

OpenStudy (agl202):

is correct answer!!!^

OpenStudy (aum):

What is said above is true. But it is good to know how to do this step-by-step too. Perimeter = 2(L + W) = 68 ---- (1) L / W = 9 / 8. So L = 9/8 * W. Replace L in (1) by 9W/8, solve for W and then solve for L.

OpenStudy (anonymous):

thank you, guys!

OpenStudy (aum):

You are welcome.